# Exact output tracking for the one-dimensional heat equation and applications to the interpolation problem in Gevrey classes of order 2

### v1

@@ 2.5 Proof of Proposition 4
In this subsection we prove Proposition 4. We will use the following Lemma, taken from [reachable_martin, Theorem 3.8].
@@
According to [reachable_martin, Remark 3.3, item 3], the function
@@
satisfies the hypotheses of Lemma 21. Hence there exists $\phi\in C^{\infty}[-1,0]$ such that
@@
where $b_{n+1}:=n!(n+1)!$ (no condition is imposed on $b_{0}$). By Stirling’s formula,
@@
$\operatorname{Li}_{s}(\zeta)=\frac{\Gamma(1-s)2^{s-1}}{(1-\zeta)^{1-s}}+O(1),\quad\zeta\rightarrow 1,\quad\zeta\in D(1/2,1/2),$
@@
We end this Section with two consequences of Proposition 4. First, we formulate below the extension problem in the class $C^{\infty}(k^{2k})$, to be compared with Definition 25.
@@
Let
$\infty<a<b<\infty,\quad A>0,\quad 1\leq\rho<\infty,$
and consider a segment $a^{\prime},c^{\prime}]$ that strictly contains the segment $[a,b]$. We say that the extension problem associated with $(a,b,a^{\prime},b^{\prime},A)$ is solvable with loss $\rho$ if: for every functions $\phi\in C^{\infty}[a,b]$ such that
$\exists C>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[a,b],\quad|\phi^{(n)}(t)|\leq CA^{n}n^{2n},$
and for every $\epsilon>0$, there exists a function $\varphi\in C^{\infty}[a^{\prime},b^{\prime}]$ such that
$\forall x\in[a,b],\quad\phi(x)=\varphi(x),$
@@
$\exists C^{\prime}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[a^{\prime},b^{\prime}],\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{2n},$
Reasoning similarly as for the proof of Proposition 26, it can be shown that the solvability of the extension problem does not depend on $(a,b,a^{\prime},b^{\prime},A)$, but only on the loss factor $\rho$. Moreover, the two following facts are almost trivial: on the one hand, if the extension problem can be solved with loss factor $\rho\geq 1$, then it can also be solved with any $\tilde{\rho}\geq\rho$. This allows to define $\rho^{Ext}$ as the optimal (i.e. minimal) loss factor $\rho$ such that the extension problem is solvable with loss $\rho$. On the other hand, if the interpolation problem (see Definition 25) can be solved with loss $\rho$, then the extension problem can be solved with the same loss $\rho$. Denoting $\rho^{Carl}$ the optimal loss factor for the interpolation problem, we thus have $\rho^{Ext}\leq\rho^{Carl}$. Proposition 4 shows that $\rho^{Ext}\geq 2$, and Mityagin’s Theorems bring $\rho^{Carl}=2$, hence $\rho^{Ext}=\rho^{Carl}$. In other words: there is no way to extend functions in Gevrey classes $C^{\infty}(k^{2k})$ which yields a better loss factor than using the solution of the interpolation problem.
@@
Consider $C^{\infty}([-1,1],k^{2k},A)$ as the set of functions $\varphi\in C^{\infty}[-1,1]$ such that $|\varphi^{(k)}(x)|\lesssim A^{k}k^{2k}$. Denote $\mathcal{B}_{x_{0}}^{A}$ the Borel operator acting on $C^{\infty}([-1,1],k^{2k},A)$, at the point $x_{0}\in[-1,1]$. We then have:
  • For every $x_{0}\in(-1,1)$, $\operatorname{Range}\mathcal{B}_{x_{0}}^{A}=\operatorname{Range}\mathcal{B}_{0}^{A}\subset\operatorname{Range}\mathcal{B}_{-1}^{A}\cap\operatorname{Range}\mathcal{B}_{1}^{A}$.
  • For every $\epsilon>0$, there holds $\operatorname{Range}\mathcal{B}_{\pm 1}^{A}\subset\operatorname{Range}\mathcal{B}_{\mp 1}^{2A+\epsilon}$.
  • For every $\epsilon>0$, we do not have $\operatorname{Range}\mathcal{B}_{\pm 1}^{A}\subset\operatorname{Range}\mathcal{B}_{\mp 1}^{2A-\epsilon}$.
@@
Indeed, the first item is trivial, the second one is by Mitjagin’s Theorem 27, and the third item is a consequence of Proposition 4. ∎
@@ 3.1.1 Neumann-to-Neumann
$\left\{\begin{array}[]{rclcc}z_{t}(t,x)&=&z_{xx}(t,x),&0<x<1,&t>0,\\ z_{x}(t,1)&=&u(t),\\ z(t,0)&=&0,\\ z(0,x)&=&0,\\ y(t)&=&z_{x}(t,0),\end{array}\right$. (37)
which we call ‘‘Neumann-to-Neumann’’ because the control is through the Neumann action and the output is the Neumann trace. When the initial data in (37) is nonzero, it is a well-posed linear and time-invariant system [LTI] on the state space $L^{2}(0,1)$. From [hartmann_orsoni, §3.2], its reachable space is
$\mathcal{R}_{\operatorname{Neu}-\operatorname{Neu}}=\{f\in\operatorname{Hol}(\Omega):f\mbox{ odd},\quad f^{\prime}\in L^{2}(\Omega)\},$
and its transfer function is
@@
$\mathcal{Y}_{\operatorname{Neu}-\operatorname{Neu}}(0,\infty)=\mathcal{G}_{2,1/\sqrt{2},0}(0,\infty)\cap C^{\infty}_{(0)}[0,\infty),$
and the trackable space in finite time is given by
$\mathcal{Y}_{\operatorname{Neu}-\operatorname{Neu}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},0}(0,T)\cap C^{\infty}_{(0)}[0,T]:\frac{\partial}{\partial\zeta}\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k+1}}{(2k+1)!}\in L^{2}(\Omega)\right\}.$
@@ 3.1.2 Dirichlet-to-Neumann
When the initial data in (38) is nonzero, it is a well-posed linear and time-invariant system on the state space $H^{-1}(0,1)$. Similarly to in the previous case, we deduce successively
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Neu}}=\{f\in\operatorname{Hol}(\Omega):f\mbox{ odd},\quad f\in L^{2}(\Omega)\},\quad\mathbf{H}_{\operatorname{Dir}-\operatorname{Neu}}(s)=\frac{\sqrt{s}}{\sinh\sqrt{s}},$
@@
$\mathcal{Y}_{\operatorname{Dir}-\operatorname{Neu}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},-1/2}(0,T)\cap C^{\infty}_{(0)}[0,T]:\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k+1}}{(2k+1)!}\in L^{2}(\Omega)\right\}.$
@@ 3.1.3 Dirichlet-to-Dirichlet
When the initial data in (39) is nonzero, it is a well-posed linear and time-invariant system on the state space $H^{-1}(0,1)$. Similarly to in the previous case, we deduce successively
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Dir}}=\{f\in\operatorname{Hol}(\Omega):f\mbox{ even},\quad f\in L^{2}(\Omega)\},\quad\mathbf{H}_{\operatorname{Dir}-\operatorname{Dir}}(s)=\frac{1}{\cosh\sqrt{s}},$
@@
$\mathcal{Y}_{\operatorname{Dir}-\operatorname{Dir}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},0}(0,T)\cap C^{\infty}_{(0)}[0,T]:\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k}}{(2k)!}\in L^{2}(\Omega)\right\}.$
@@ 3.2 Smooth control laws
Consider the case where the control $u$ belongs to $H^{p}_{(0)}(0,\infty)$, the set of functions such that $u,u^{\prime},...,u^{(p)}$ every belong to $L^{2}(0,\infty)$ and for which $u(0)=...=u^{(p-1)}(0)=0$. For such $u$, the relation (12) can be rewritten using
@@
This allows to identify the set $\mathcal{Y}^{p}(0,\infty)$ of these signals that can be tracked with $H^{p}_{(0)}(0,\infty)$ controls. For finite time considerations, Theorem 2 can be adapted using the knowledge we have of the reachable space of the heat equation with smooth inputs, see [hartmann_orsoni, §3.2]. In fact one can alternatively use the following fact, more general and not relying on any peculiar structure. Consider a well-posed LTI system $\Sigma(A,B,C)$ with input space $U$ and outputs denoted by $y$. As a consequence, a signal $y$ is an output of the system with control lying in $H^{p}_{(0)}(0,T;U)$ if and only if $y,y^{\prime},...,y^{(p)}$ are every outputs of the system with controls $L^{2}(0,T;U)$.
@@ 3.3 Other point measurement
$\left\{\begin{array}[]{rclcc}z_{t}(t,x)&=&z_{xx}(t,x),&0<x<1,&t>0,\\ z_{x}(t,1)&=&u(t),\\ z_{x}(t,0)&=&0,\\ z(0,x)&=&0,\\ y(t)&=&z(t,x_{0}),\end{array}\right$. (40)
@@
The new Gevrey radius can be easily guessed: the number $(1-x_{0})/\sqrt{2}$ is the radius of the circle centered at $x_{0}$ and inscribed in $\Omega$.
@@
Let $u\in L^{2}(0,\infty)$ and denote $y$ the associated output of (40). Applying the Laplace transform, we obtain
$\hat{y}(s)=\frac{\cosh(\sqrt{s}x_{0})}{\sqrt{s}\sinh(\sqrt{s})}\hat{u}(s),\quad\operatorname{Re}s>0,$
which we rewrite as
@@
Conversely, let $\varphi\in C^{\infty}_{(0)}[0,\infty)$ satisfy (41). Clearly it is enough to show that $\psi\hat{\dot{y}}\in H^{2}(\mathbb{C}_{+})$. We use Proposition 17, the latter function has boundary values $\psi(i\xi)\mathcal{F}\dot{y}(\xi)$ which are square- integrable by hypothesis (41) and Theorem 7. It is therefore enough to check the condition on the support of the Fourier transform. To simplify the notations, we introduce the shorthand $\operatorname{spec}T:=\operatorname{supp}\mathcal{F}T$ for every $T\in\mathcal{S}^{\prime}(\mathbb{R})$. We then have
@@
4 Interpolation in Gevrey classes
In this Section we prove Proposition 10. The proof of Corollary 9 is straightforward and omitted.

### v2

@@ 2.5 Proof of Proposition 4
In this subsection we prove Proposition 4. We will rely on the following Lemma taken from [reachable_martin, Theorem 3.8], which is proved using the Borel summation [ramis].
@@
Following [reachable_martin, Remark 3.3, item 3], we consider the function
@@
which satisfies the hypotheses of Lemma 24. Hence there exists $\phi\in C^{\infty}[-1,0]$ such that
@@
where $b_{n+1}:=n!(n+1)!$ and $b_{0}=1$. By Stirling’s formula,
@@
$\operatorname{Li}_{s}(\zeta)=\frac{\Gamma(1-s)2^{s-1}}{(1-\zeta)^{1-s}}+O(1),\quad\zeta\rightarrow 1,\quad\zeta\in D(1/2,1/2),$ (38)
@@
We end this Section with two consequences of Proposition 4. First, we formulate below the extension problem for Gevrey functions of order $2$, to be compared with Definition 28.
@@
We say that the extension problem (for $s=2$) is solvable with loss $\rho\in(0,1)$ if: for every functions $\phi\in C^{\infty}[-1,1]$ such that
$\exists C,R_{0}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[-1,1],\quad|\phi^{(n)}(t)|\leq C\frac{(2n)!}{R_{0}^{2n}},$
and for every $\epsilon>0$, there exists a function $\varphi\in C^{\infty}[-2,2]$ such that
$\forall t\in[-1,1],\quad\phi(t)=\varphi(t),$
@@
$\exists C^{\prime}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[-2,2],\quad|\varphi^{(n)}(t)|\leq C^{\prime}\frac{(2n)!}{(\rho R_{0}-\epsilon)^{2n}}.$
It is clear that if the extension problem can be solved with some loss factor $\rho$, then it can also be solved with any $\tilde{\rho}\leq\rho$. This allows to define $\rho^{Ext}$ as the optimal (i.e. maximal) loss factor $\rho$ such that the extension problem is solvable with loss $\rho$. Recall that from Mitjagin’s theorem the interpolation problem can be solved for Gevrey functions of order $2$, the optimal loss factor being $\rho^{Carl}=1/\sqrt{2}$. Proposition 4 shows that $\rho^{Ext}\leq 1/\sqrt{2}$ and it easy to see that $\rho^{Ext}\geq\rho^{Carl}$, hence $\rho^{Ext}=\rho^{Carl}$. In other words: there is no way to extend functions in the Gevrey class of order $2$ which yields a better loss factor than using the solution of the interpolation problem.
@@
Let $R>0$ and denote $\mathcal{B}_{x_{0}}^{R}$ the Borel operator acting on $G^{2,R}[-1,1]$, at the point $x_{0}\in[-1,1]$. We then have:
  • For every $x_{0}\in(-1,1)$, $\operatorname{Range}\mathcal{B}_{x_{0}}^{R}=\operatorname{Range}\mathcal{B}_{0}^{R}\subset\operatorname{Range}\mathcal{B}_{-1}^{R}\cap\operatorname{Range}\mathcal{B}_{1}^{R}$.
  • For every $\epsilon>0$, there holds $\operatorname{Range}\mathcal{B}_{\pm 1}^{R}\subset\operatorname{Range}\mathcal{B}_{\mp 1}^{R/\sqrt{2}-\epsilon}$.
  • For every $\epsilon>0$, we do not have $\operatorname{Range}\mathcal{B}_{\pm 1}^{R}\subset\operatorname{Range}\mathcal{B}_{\mp 1}^{R/\sqrt{2}+\epsilon}$.
@@
The first item is trivial, the second one is by Mitjagin’s Theorem 30, and the third item is a consequence of Proposition 4. ∎
@@ 3.1.1 Neumann-to-Neumann
$\left\{\begin{array}[]{rclcc}z_{t}(t,x)&=&z_{xx}(t,x),&0<x<1,&t>0,\\ z_{x}(t,1)&=&u(t),\\ z(t,0)&=&0,\\ z(0,x)&=&z^{0}(x),\\ y(t)&=&z_{x}(t,0),\end{array}\right$. (39)
which we call “Neumann-to-Neumann” because the control is through the Neumann action and the output is the Neumann trace. Let us first disregard the output $y$; similarly as for (1) we consider transposition solutions and it is easy to check that (39) is well-posed on the state space $L^{2}(0,1)$. When $z^{0}=0$ and $u\in L^{2}(0,\infty)$, extending the solution $z$ of (39) as an odd function of $x\in(-1,1)$ and using the hypoellipticity of the heat equation one sees that $z\in C^{\infty}([0,\infty)\times(-1,1))$. Hence, when $z^{0}=0$, the output $y(t)=z_{x}(t,0)$ is just a pointwise evaluation. When $u=0$ and $z^{0}\in L^{2}(0,1)$, one sees that $z\in C^{\infty}((0,\infty)\times(-1,1))$ and $z_{x}(\cdot,0)\in L^{1}(0,\infty)$, but we will not use this fact.
Let now $u\in L^{2}(0,\infty)$, we compute the Laplace transform of $y$. A representation formula similar to (33) would yield $z\in C([0,\infty)\times[0,1])$ but be of no use here because the output is the Neumann trace, rather we shall pass the equation (39) to the Laplace transform. To this aim, we first take $u\in H^{1}(0,\infty)$ such that $u(0)=0$. We consider the function
$\xi(t,x)=z(t,x)-x(x-1)u(t),$
which is continuous in both variables and solves
$\left\{\begin{array}[]{rclcc}\xi_{t}(t,x)&=&\xi_{xx}(t,x)+f(t,x),&0<x<1,&t>0,\\ \xi_{x}(t,1)&=&0,\\ \xi(t,0)&=&0,\\ \xi(0,x)&=&0,\end{array}\right.,\quad f(t,x)=-x(x-1)\dot{u}(t)+\frac{u(t)}{2}.$
The source term $f$ is of class $L^{2}(0,\infty;L^{2}(0,1))$, hence by maximal regularity [bensoussan_control, Proposition 3.7, §II.1] we have
$\xi\in H^{1}(0,\infty;L^{2}(0,1))\cap L^{2}(0,\infty;H^{2}(0,1)).$
The function $z$ is therefore also in the above class and we are allowed to pass (39) to the Laplace transform with respect to $t$: for all $s\in\mathbb{C}_{+}$, the function $\hat{z}(s,\cdot)$ satisfies
$\left\{\begin{array}[]{rcl c}s\hat{z}(s,x)&=&\partial_{xx}\hat{z}(s,x),&\mbox{a.e. }x\in(0,1),\\ \partial_{x}\hat{z}(s,1)&=&\hat{u}(s),\\ \hat{z}(s,0)&=&0,\end{array}\right$.
from which we deduce that
$\hat{z}(s,x)=\frac{\hat{u}(s)\sinh(\sqrt{s}x)}{\sqrt{s}\cosh(\sqrt{s})}.$
We then evaluate
$\hat{y}(s)=\partial_{x}\hat{z}(s,0)=\frac{\hat{u}(s)}{\cosh\sqrt{s}},$
so that the transfer function of (39) is thus
@@
Notice that the above function is bounded in modulus over $\mathbb{C}_{+}$, from which one deduces by approximation arguments that for general $u\in L^{2}(0,\infty)$, the output $y$ is $L^{2}(0,\infty)$.
@@
$\mathcal{Y}_{\operatorname{Neu}-\operatorname{Neu}}(0,\infty)=\mathcal{G}_{2,1/\sqrt{2},0}(0,\infty)\cap C^{\infty}_{(0)}[0,\infty).$
For the trackable subspace in finite time, we need the reachable space of the control system (39), that is defined by the same equations as (39) but with nonzero initial data disregarding the output $y$. From [hartmann_orsoni, §3.2], this reachable space is given by
$\mathcal{R}_{\operatorname{Neu}-\operatorname{Neu}}=\{f\in\mathcal{H}(\Omega):f\mbox{ odd},\quad f^{\prime}\in L^{2}(\Omega)\}.$
Repeating the proof of Theorem 2, one arrives to
$\mathcal{Y}_{\operatorname{Neu}-\operatorname{Neu}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},0}(0,T)\cap C^{\infty}_{(0)}[0,T]:\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k}}{(2k)!}\in A^{2}(\Omega)\right\}.$
@@ 3.1.2 Dirichlet-to-Neumann
When the initial data in (40) is nonzero and disregarding the output $y$, it is a well-posed system on the state space $H^{-1}(0,1):=(H^{1}_{0}(0,1))^{\prime}$, with reachable space
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Neu}}=\{f\in\mathcal{H}(\Omega):f\mbox{ odd},\quad f\in L^{2}(\Omega)\}.$
Observe that when the initial data $z^{0}\in H^{-1}(0,1)$ of (40) is nonzero and the control $u$ is zero, the output $y(t)=z_{x}(t,0)$ may fail to be integrable around $t=0$. But this does not affect our method as explained in Remark 23. Reasoning as in the previous subsubsection one sees that the transfer function of (40) is given by
$\mathbf{H}_{\operatorname{Dir}-\operatorname{Neu}}(s)=\frac{\sqrt{s}}{\sinh\sqrt{s}},$
@@
$\mathcal{Y}_{\operatorname{Dir}-\operatorname{Neu}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},-1/2}(0,T)\cap C^{\infty}_{(0)}[0,T]:\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k+1}}{(2k+1)!}\in A^{2}(\Omega)\right\}.$
@@ 3.1.3 Dirichlet-to-Dirichlet
When the initial data in (41) is nonzero and disregarding the output $y$, it is a well-posed system on the state space $H^{-1}(0,1)$. We have
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Dir}}=\{f\in\mathcal{H}(\Omega):f\mbox{ even},\quad f\in L^{2}(\Omega)\},\quad\mathbf{H}_{\operatorname{Dir}-\operatorname{Dir}}(s)=\frac{1}{\cosh\sqrt{s}},$
@@
$\mathcal{Y}_{\operatorname{Dir}-\operatorname{Dir}}(0,T)=\left\{y\in\mathcal{G}_{2,1/\sqrt{2},0}(0,T)\cap C^{\infty}_{(0)}[0,T]:\sum_{k=0}^{\infty}y^{(k)}(T)\frac{\zeta^{2k}}{(2k)!}\in A^{2}(\Omega)\right\}.$
@@ 3.2 Smooth control laws
For a given integer $p\geq 1$, consider the case where the control $u$ belongs to $H^{p}_{0}(0,\infty)$, which is defined as
$H^{p}_{0}(0,\infty):=\left\{u\in H^{p}(0,\infty):u(0)=u^{\prime}(0)=\cdots=u^{(p-1)}(0)=0\right\}.$
For such $u$, the relation (12) can be rewritten using
@@
This allows to identify the set $\mathcal{Y}^{p}(0,\infty)$ of these signals that can be tracked with $H^{p}_{0}(0,\infty)$ controls. For finite time considerations, Theorem 2 can be adapted using the knowledge we have of the reachable space of the heat equation with smooth inputs, see [hartmann_orsoni, §3.2].
@@ 3.3 Other point measurement
$\left\{\begin{array}[]{rclcc}z_{t}(t,x)&=&z_{xx}(t,x),&0<x<1,&t>0,\\ z_{x}(t,1)&=&u(t),\\ z_{x}(t,0)&=&0,\\ z(0,x)&=&0,\\ y(t)&=&z(t,x_{0}).\end{array}\right$. (42)
@@
The Gevrey radius $(1-x_{0})/\sqrt{2}$ is the radius of the circle centered at $x_{0}$ and inscribed in $\Omega$.
@@
Let $u\in L^{2}(0,\infty)$ and denote $y$ the associated output of (42). Reasoning as in the proof of Theorem 1555One could also reason as for (39) considering the output $z_{t}(t,x_{0})$ we deduce that $y$ is Laplace transformable, with absicssa of absolute convergence no greater than 0, and satisfies
$\hat{y}(s)=\frac{\cosh(\sqrt{s}x_{0})}{\sqrt{s}\sinh(\sqrt{s})}\hat{u}(s),\quad\operatorname{Re}s>0.$
We rewrites the above as
@@
Conversely, let $\varphi\in C^{\infty}_{(0)}[0,\infty)$ satisfy (43). Clearly it is enough to show that $\psi\hat{\dot{y}}\in\mathcal{H}^{2}(\mathbb{C}_{+})$. We use Proposition 18, the latter function has boundary values $\psi(i\xi)\mathcal{F}\dot{y}(\xi)$ which are square- integrable by hypothesis (43) and Theorem 7. It is therefore enough to check the condition on the support of the Fourier transform. To simplify the notations, we introduce the shorthand $\operatorname{spec}T:=\operatorname{supp}\mathcal{F}T$ for every $T\in\mathcal{S}^{\prime}(\mathbb{R})$. We then have
@@
4 Interpolation in Gevrey classes of order 22
In this Section we prove Proposition 9 and Corollary 10.